Grade 12Algebra

Find
3.96 ÷ 4
Find
3.96 ÷ 4
Find
3.96 ÷ 4
Find
3.96 ÷ 4

Pawan Prajapati
5 Years agoGrade 12

1 Answer

Harshit Singh
5 Years ago
Dear Student

We have,
= (396/100) ÷ 4
Then,
= (396/100) × (1/4)
= (99/100) × (1/1)
= (99/100)
On dividing a decimal by 100, the decimal point is shifted to the left by two places.
Then,
= 0.99

Thanks